Consider a project network that begins with activity A and is followed by activity B. Activity A takes 6 days to finish. If the earliest start time for activity B is 6 days and the latest start time for it is 10 days, activity B has a slack of _____. 4 days 16 days 10 days 12 days
Answer #1

ANSWER:

Slack time = latest start time of activity b - earliest start time of activity b = 10 - 6 = 4

so the correct answer is option a that is 4 days.
Answer #1

ANSWER:

Slack time = latest start time of activity b - earliest start time of activity b = 10 - 6 = 4

so the correct answer is option a that is 4 days.